Actions
Ops/Comm Scheduling meeting : 24/10/2017¶
Location : CR21 at 14:00 pm
Phone access : phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517
- MOM update
- Data quality
- Detectors : Durga Run10079_3-170.pdf 3-vs-10-170.pdf
- Physics shifter
- Target : Joe
- Target replacement issues : Henry
- Shutdown activities - the list so far
Name | Task | Time required |
Henry | Restart conventional magnet controls crate | 1-2 hours (?) |
Josef | Replace filters in 8 compressors on mezzanine | 1-2 days |
Ajit | Test AMI controllers | 0.5 - 1 day |
Mark | Partial warming of DS to clean vacuum space and recool | 10 days |
Ajit/MOM | Review current alarms | 1 day |
Joe/Henry | Change target patch fibres | 0.5 days |
John/Colin | Hall AC work? | ? |
- Others?
- Run Plan (not sure how current this now): 2017-09-28-Detailed-RunPlan-v7.xlsx
- Plans for next user cycle : 2017/03 : Nov 14 - Dec 20
- We have permission to run in the next user cycle.
- Plan is to take data with empty LH2 absorber, and with LiH absorber with M2 (plus other measurements if time and manpower is available)
- Shifts are now open. Please sign up. 16/48 shifts have been assigned.
- MOMs will be : Durga, Victoria+Melissa, Paolo
- Run up will probably be on Nov 14th with data-taking starting on Nov 15th.
- AOB
Updated by Franchini, Paolo almost 6 years ago · 10 revisions